# Quest for optimal quantum resetting: protocols for a particle on a chain

### v3

@@ IV Results
We will now try to explain the above-mentioned nature of $S^{r}_{n}$ for different $p$ values. The probability distribution of $j^{*}$ (i.e., the site where the particle gets resets each time) is completely analogous to the probability distribution of the position of a one-dimensional biased discrete random walk (see Eq. (LABEL:eqn_6)). The interval $\Delta$ and the number of resetting steps $R$ in our strategies are analogous to the step length and the number of total jumps associated with a random walk problem. In the random-walk picture, the $p$ ($1-p$) value can be thought of as the probability of going to the right (left) sites. Drawing the mapping between a random walk problem and MPR protocol can help us to compute the mean and variance of the $P(j^{*})$, which are given by,
@@ V Discussion
For a classical particle, doing an unbiased random walk on a one-dimensional lattice where at a given site a detector is placed, it is well established that the average hitting time or mean FDT reduces if one resets the particle on its original position again and again with an optimal time interval. Interestingly, if one computes the probability distribution of the sites visited by that classical stochastic particle, the mean and peak of the distribution both lie in the same position and that coincides with the particle’s initial position as well. However, this is not the case for a quantum particle moving on a tight-binding lattice. As a matter of fact, it turns out that the mean of the distribution coincides with the initial position of the quantum particle, but the peak does not. On the other hand, the sites where the distribution is peaked signify where the particle is most likely to be found. According to our resetting protocols, we reset the particle at those sites with some associated probabilities to the right and left (as there is a two-fold degeneracy in those sites where the probability of finding the particle is maximum) instead of its initial position and found that the efficiency of the search process increased significantly compared to the case where resetting is done at the initial position, with guaranteed detection when the associated probabilities to the right peak and the left peak are equal. We believe that our MPR resetting protocol can be a natural choice in the quantum context that has no classical analog. We would like to point out that we had restricted ourselves in this work to the nearest-neighbor hopping model on a 1D chain, which is exactly solvable; hence, finding the most likely position is straightforward. However, even if we do not have a solvable system, one can easily numerically study the unitary evolution to find out the most probable positions as a function of $t_{r}$ for such models. Hence, in principle, our study can be extended even for non-solvable models with a bit of extra effort.
@@ Appendix B $\tau$ dependence on Mean FDT
Appendix C Most probable positions under repetitive projective measurements
@@
Figure. 11 shows the time evolution of the norm square of the wave function starting from the site $j=0$ differs quite a lot when the wave function evolves under measurement-free unitary dynamics compared to a situation where the dynamics get interrupted by projective measurements at site $\delta=10$. In both cases, the most probable positions are not the same. While according to the Unitary dynamics, the peaks of the distribution should be situated at the sites $j=10$ and $j=-10$ at time $t=6$, due to the repetitive measurements, the wave function profile has changed, and the peak of the distribution does not remain as at $j=10$. Hence, potentially the path-dependent protocols could be harder to implement; we restrict ourselves to the path-independent MPR protocol in the main text, where the knowledge of resetting history is not required.

### v4

@@ IV Results
We will now try to explain the above-mentioned nature of $S^{r}_{n}$ for different $p$ values. The probability distribution of $j^{*}$ (i.e., the set of possible resetting positions at a given resetting step) is completely analogous to the probability distribution of the position of a one-dimensional biased discrete random walk (see Eq. (LABEL:eqn_6)). The interval $\Delta$ and the number of resetting steps $R$ in our strategies are analogous to the step length and the number of total jumps associated with a random walk problem. In the random-walk picture, the $p$ ($1-p$) value can be thought of as the probability of going to the right (left) sites. Drawing the mapping between a random walk problem and MPR protocol can help us to compute the mean and variance of the $P(j^{*})$, which are given by,
@@ V Discussion
For a classical particle, doing an unbiased random walk on a one-dimensional lattice where at a given site a detector is placed, it is well established that the average hitting time or mean FDT reduces if one resets the particle on its original position again and again with an optimal time interval. Interestingly, if one computes the probability distribution of the sites visited by that classical stochastic particle, the mean and peak of the distribution both lie in the same position and that coincides with the particle’s initial position as well. However, this is not the case for a quantum particle moving on a tight-binding lattice. As a matter of fact, it turns out that the mean of the distribution coincides with the initial position of the quantum particle, but the peak does not. On the other hand, the sites where the distribution is peaked signify where the particle is most likely to be found. According to MPR protocol, at a given resetting step, resets are done with certain probabilities to the set of possible peak positions that could occur because of the previous resets and followed by uninterrupted unitary evolution, irrespective of which path was taken by the particle in previous steps. We have found that the efficiency of the search process increased significantly compared to the case where resetting is done at the initial position, with guaranteed detection when the associated probabilities of going to the right and the left at each random walk step required to reach a possible resetting position($R$ number of random walk steps at the $R$-th resetting step) are equal. We believe that our MPR resetting protocol can be a natural choice in the quantum context that has no classical analog. We would like to point out that we had restricted ourselves in this work to the nearest-neighbor hopping model on a 1D chain, which is exactly solvable; hence, finding the most likely position is straightforward. However, if we do not have a solvable system, one can numerically study the unitary evolution to find out the most probable positions as a function of $t_{r}$ for such models. Hence, in principle, our study can be extended even for non-solvable models with extra classical simulations. In that case, more rigorous and system-dependent studies are required to assure quantum advantage.
@@ Appendix B $\tau$ dependence on Mean FDT
Appendix C Evolution under repetitive projective measurements
@@
Figure. 11 shows the time evolution of the norm square of the wave function starting from the site $j=0$ interrupted by projective measurements at site $\delta=10$.